# Dual-Head Knowledge Distillation: Enhancing Logits Utilization with an Auxiliary Head

### v1

@@ 3.3 Our Proposed Method
Figure 4: An illustration of Dual-Head Knowledge Distillation (DHKD). DHKD decouples the original linear classifier into a duo. We can evade conflicts between two losses by introducing an Adaptive Auxiliary Classifier customized according to the models’ architectures. The right side shows the gradient alignment method on CIFAR-100. When the angle between two gradients is larger than $90^{\circ}$, we will project the gradient of the BinaryKL loss to the orthogonal direction of the gradient of the CE loss.
For a given input $\bm{x}$, we can get a feature $\bm{h}=f(\bm{x})$ through the backbone $f(\cdot)$. Just as Figure 4 shows, we use two classifiers for different losses: $g(\cdot)$ for the CE loss and $g^{\prime}(\cdot)$ for the BinaryKL loss.
The original BinaryKL loss suffers from a problem: when the student output falls into a small neighborhood of the teacher output (which means it is close to the optimization goal), the derivative of the BinaryKL loss at the student output also depends on the value of the teacher output. As a result, the optimization progress cannot be synchronized among distinct teacher outputs with the same $\tau$. Although using different $\tau$ in different settings can lead to state-of-the-art performance, it would introduce more hyper-parameters. To mitigate this issue, we use a variant of the BinaryKL loss. To unify the gradients at the outputs of the teacher model, we choose to narrow the distance between zero and the difference between the teacher and student. The modified loss (we call it BinaryKL-Norm) can be formulated as follows:
$\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}=\tau^{2}\sum\limits_{i=1}^{B}\sum\limits_{k=1}^{K}\mathcal{KL}\left(\left.\left[\frac{1}{2},\frac{1}{2}\right]\right\|\left[\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right),1-\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right)\right]\right).$ (5)
@@
$\mathcal{G}_{\mathrm{BKL}}^{\mathrm{proj}}=\begin{cases}\mathcal{G}_{\mathrm{BKL}},&\quad\text{if}\ \mathcal{G}_{\mathrm{BKL}}\cdot\mathcal{G}_{\mathrm{CE}}\geq 0,\\ \mathcal{G}_{\mathrm{BKL}}-\frac{\mathcal{G}_{\mathrm{BKL}}\cdot\mathcal{G}_{\mathrm{CE}}}{\|\mathcal{G}_{\mathrm{CE}}\|^{2}}\mathcal{G}_{\mathrm{CE}},&\quad\text{otherwise}.\\ \end{cases}$ (6)
@@
Table 1: Results on the CIFAR-100 validation. Teachers and students are in the same architectures. The best performance is highlighted in bold, and the second best performance is highlighted with an underline.
Teacher resnet56 resnet110 resnet32$\times$4 WRN-40-2 WRN-40-2 VGG13
72.34 74.31 79.42 75.61 75.61 74.64
Student resnet20 resnet32 resnet8$\times$4 WRN-16-2 WRN-40-1 VGG8
69.06 71.14 72.50 73.26 71.98 70.36
features FitNet 69.21 71.06 73.50 73.58 72.24 71.02
RKD 69.61 71.82 71.90 73.35 72.22 71.48
CRD 71.16 73.48 75.51 75.48 74.14 73.94
OFD 70.98 73.23 74.95 75.24 74.33 73.95
ReviewKD 71.89 73.89 75.63 76.12 75.09 $\bm{74.84}$
logits KD 70.66 73.08 73.33 74.92 73.54 72.98
DKD $\bm{71.97}$ $\bm{74.11}$ 76.32 76.24 74.81 74.68
DHKD 71.19 73.92 $\bm{76.54}$ $\bm{76.36}$ $\bm{75.25}$ $\bm{74.84}$
Table 2: Results on the CIFAR-100 validation. Teachers and students are in different architectures. The best performance is highlighted in bold, and the second best performance is highlighted with an underline.
Teacher resnet32$\times$4 WRN-40-2 VGG13 ResNet-50 resnet32$\times$4
79.42 75.61 74.64 79.34 79.42
Student ShuffleNet-V1 ShuffleNet-V1 MBN-V2 MBN-V2 ShuffleNet-V2
70.50 70.50 64.60 64.60 71.82
features FitNet 73.59 73.73 64.14 63.16 73.54
RKD 72.28 72.21 64.52 64.43 73.21
CRD 75.11 76.05 69.73 69.11 75.65
OFD 75.98 75.85 69.48 69.04 76.82
ReviewKD $\bm{77.45}$ 77.14 $\bm{70.37}$ 69.89 77.78
logits KD 74.07 74.83 67.37 67.35 74.45
DKD 76.45 76.70 69.71 70.35 77.07
DHKD 76.78 $\bm{77.25}$ 70.09 $\bm{71.08}$ $\bm{77.99}$
4 Experiments
Information about the comparing methods and implementation details can be found in Appendices E and F. More experimental results and visualizations can be found in Appendices A, G, H, I, J and K.
4.1 Main Results
CIFAR-100 image classification. The validation accuracy on CIFAR-100 is reported in Table 1 and Table 2. Table 1 shows the results where teachers and students are of the same network architectures. Table 2 contains the results where teachers and students are from different architectures. It can be observed that our DHKD achieves remarkable improvements over the vanilla KD on all teacher-student pairs. Compared with the state-of-the-art logit-based method, our method can achieve better performance except in two experiments across ResNet architectures. Furthermore, DHKD achieves comparable or even better performance than feature-based methods, significantly improving the trade-off between performance and training efficiency.
ImageNet image classification. The top-1 and top-5 validation accuracy on ImageNet is reported in Table 3 and Table 4. Table 3 shows the results where teachers and students have the same network architectures. Table 4 shows the results where teachers and students have different architectures. We can find that our DHKD achieves comparable or even better performance than the existing methods. The success on the large-scale dataset further proves the effectiveness of our method.
4.2 Ablation Studies
@@
Table 3: Top-1 and top-5 accuracy (%) on the ImageNet validation. We set ResNet-34 as the teacher and ResNet-18 as the student.
distillation manner features logits
Metric Teacher Student AT OFD CRD ReviewKD KD DKD DIST DHKD
top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.07 $\bm{72.15}$
top-5 91.42 89.07 90.01 89.98 90.13 90.51 89.88 90.41 90.42 $\bm{90.89}$
Table 4: Top-1 and top-5 accuracy (%) on the ImageNet validation. We set ResNet-50 as the teacher and MobileNet as the student.
distillation manner features logits
Metric Teacher Student AT OFD CRD ReviewKD KD DKD DIST DHKD
top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 $\bm{73.24}$ 72.99
top-5 92.86 88.76 89.33 90.34 90.41 91.00 88.98 91.05 91.12 $\bm{91.45}$
Table 5: Ablation studies on CIFAR-100. We choose one pair of models with the same architecture and another pair with different architectures. For the former pair, we set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student; for the latter pair, we set ResNet-50 as the teacher and MobileNet-V2 as the student.
CE BinaryKL -Norm Dual Head Gradient Alignment Nonlinear Auxiliary Head resnet32$\times$4 resnet8$\times$4 ResNet-50 MBN-V2
72.50 64.60
75.15 68.52
N.A. N.A.
76.16 70.16
$\bm{76.54}$ 69.97
74.38 $\bm{71.08}$
5 Conclusion
@@ Appendix A Dual Head is the Only Way
Table 6: Some attempts that change the learning rate. Top-1 accuracy on CIFAR-100 is given in the table. For elements that are “N.A.”, we give the ordinal number of the epoch when the models collapse during the training phase. It can be observed that almost all the student models collapse. As the learning rate decreases, the time for model collapse is delayed.
Same Architecture Different Architectures
teacher resnet56 resnet110 resnet32$\times$4 resnet32$\times$4 VGG13 ResNet-50
student resnet20 resnet32 resnet8$\times$4 ShuffleNet-V1 MBN-V2 MBN-V2
lr=1e-2 N.A.(8) N.A.(6) N.A.(5) N.A.(8) N.A.(44) N.A.(34)
lr=5e-3 N.A.(8) N.A.(8) N.A.(11) N.A.(16) N.A.(48) N.A.(69)
lr=1e-3 N.A.(35) N.A.(29) N.A.(27) N.A.(50) N.A.(126) 39.56
Table 7: Some attempts using gradient clipping (GC) method. We change the Maximum Gradient Norm Value (MGNV) in a small step size, and the student model still cannot achieve comparable performance with the model trained with only the CE loss. We set resnet56 as the teacher and resnet20 as the student. All the experiments are done on CIFAR-100.
MGNV 0.01 0.1 0.3 0.5 0.7 1.0 Pure CE No GC
Test Accuracy 1.95 23.63 59.11 65.16 N.A. N.A. 69.06
@@ Appendix B Supplementary Material about Neural Collapse
B.1 Detailed Definition of ETF and Neural Collapse
Definition 1 (Simplex Equiangular Tight Frame (Papyan et al., 2020))
@@ B.1 Detailed Definition of ETF and Neural Collapse
$\mathbf{M}=\sqrt{\frac{K}{K-1}}\mathbf{U}\left(\mathbf{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right),$ (7)
@@
$\mathbf{m}_{i}^{\top}\mathbf{m}_{j}=\frac{K}{K-1}\delta_{i,j}-\frac{1}{K-1},\forall i,j\in[1,K],$ (8)
@@
B.2 Proof of Propositions
B.2.1 Proof of Proposition 1
@@
$-\frac{\partial\mathcal{L}_{CE}}{\partial\bm{w}_{k}}=\underbrace{\sum_{i=1}^{n_{k}}\left(1-p_{k}\left(\bm{h}^{\mathcal{S}}_{k,i}\right)\right)\bm{h}^{\mathcal{S}}_{k,i}}_{\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}}+\underbrace{\left(-\sum_{k^{\prime}\neq k}^{K}\sum_{j=1}^{n_{k^{\prime}}}p_{k}\left(\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)}_{\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}},$ (9)
@@
$-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL}}}{\partial\bm{w}_{k}}=\underbrace{\sum_{i=1}^{n_{k}}\tau(q_{k}(\bm{h}^{\mathcal{T}}_{k,i})-q_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i}}_{\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}}+\underbrace{\left(-\sum_{k^{\prime}\neq k}^{K}\sum_{j=1}^{n_{k^{\prime}}}\tau(q_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j})-q_{k}(\bm{h}^{\mathcal{T}}_{k^{\prime},j}))\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)}_{\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}},$ (11)
@@
$q_{k}(\bm{h})=\frac{1}{1+e^{-\bm{h}^{\top}\bm{w}_{k}/\tau}}.$ (12)
Based upon Equation 9, 11, we can prove Proposition 1.
B.2.2 Proof of Proposition 2
The definitions of $p_{k}(\bm{h})$ and $q_{k}(\bm{h})$ are the same as Equation 10 and 12.
@@ B.2.2 Proof of Proposition 2
$-\frac{\partial\mathcal{L}_{CE}}{\partial\bm{h}}=\underbrace{(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}}+\underbrace{(-\sum\limits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k})}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}}.$ (13)
@@
$-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL}}}{\partial\bm{w}_{k}}=\underbrace{\tau(q_{c}(\bm{h}^{\mathcal{T}})-q_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}}+\underbrace{\left(-\tau\sum\limits_{k\neq c}^{K}(p_{k}(\bm{h}^{\mathcal{S}})-p_{k}(\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{k}\right)}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}}.$ (14)
Based upon Equation 13, 14, we can prove Proposition 2.
B.3 Propositions for the Modified Loss

### v2

@@ 3.3. Our Proposed Method
For a given input $\bm{x}$, we can get a feature $\bm{h}=f(\bm{x})$ through the backbone $f(\cdot)$. Just as Figure 4 shows, we use two classifiers for different losses: $g(\cdot)$ for the CE loss and $g^{\prime}(\cdot)$ for the BinaryKL loss. The original BinaryKL loss suffers from a problem: when the student output falls into a small neighborhood of the teacher output (which means it is close to the optimization goal), the derivative of the BinaryKL loss at the student output also depends on the value of the teacher output. As a result, the optimization progress cannot be synchronized among distinct teacher outputs with the same $\tau$. Although using different $\tau$ in different settings can lead to state-of-the-art performance, it would introduce more hyper-parameters. To mitigate this issue, we use a variant of the BinaryKL loss. To unify the gradients at the outputs of the teacher model, we choose to narrow the distance between zero and the difference between the teacher and student. The modified loss (we call it BinaryKL-Norm) can be formulated as follows:
$\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}=\tau^{2}\sum\nolimits_{i=1}^{B}\sum\nolimits_{k=1}^{K}$ $\mathcal{KL}\Big{(}\left.\left[\frac{1}{2},\frac{1}{2}\right]\right\|$
(7) $\Big{[}\sigma\left(\left(z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}\right)/\tau\right),1-$ $\sigma\left(\left(z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}\right)/\tau\right)\Big{]}\Big{)}.$
@@
(8) $\mathcal{G}_{\mathrm{BKL}}^{\mathrm{proj}}=\begin{cases}\mathcal{G}_{\mathrm{BKL}},&\quad\text{if}\ \mathcal{G}_{\mathrm{BKL}}\cdot\mathcal{G}_{\mathrm{CE}}\geq 0,\\ \mathcal{G}_{\mathrm{BKL}}-\frac{\mathcal{G}_{\mathrm{BKL}}\cdot\mathcal{G}_{\mathrm{CE}}}{\|\mathcal{G}_{\mathrm{CE}}\|^{2}}\mathcal{G}_{\mathrm{CE}},&\quad\text{otherwise}.\\ \end{cases}$
@@
Table 2. Results on the CIFAR-100 validation. Teachers and students are in different architectures. The best performance of a single method is highlighted with a star , and the best logits-based performance is highlighted in bold. The combination of DHKD and ReviewKD always achieves the best performance and is highlighted with gray backgrounds.
[…]
. Teacher resnet32$\times$4 WRN-40-2 VGG13 ResNet-50 resnet32$\times$4 79.42 75.61 74.64 79.34 79.42 Student ShuffleNet-V1 ShuffleNet-V1 MBN-V2 MBN-V2 ShuffleNet-V2 70.50 70.50 64.60 64.60 71.82 features FitNet 73.59 73.73 64.14 63.16 73.54 RKD 72.28 72.21 64.52 64.43 73.21 CRD 75.11 76.05 69.73 69.11 75.65 OFD 75.98 75.85 69.48 69.04 76.82 ReviewKD 77.45 77.14 70.37 69.89 77.78 SimKD 77.18 77.23 69.45 71.12 78.39 CAT-KD 78.26 77.35 69.13 71.36 78.41 logits KD 74.07 74.83 67.37 67.35 74.45 DKD 76.45 76.70 69.71 70.35 77.07 DHKD 76.78 77.25 70.09 71.08 77.99 DHKD + ReviewKD 78.35 78.22 71.01 71.51 78.83
4. Experiments
Information about the comparing methods and implementation details can be found in Appendices E and F. More experimental results can be found in Appendices A, G, H and I.
4.1. Main Results
CIFAR-100 image classification.
The validation accuracy on CIFAR-100 is reported in Table 1 for cases where teachers and students share the same architecture and in Table 2 for cases where they have different architectures. It can be observed that DHKD achieves remarkable improvements over the vanilla KD on all teacher-student pairs. Compared with the SOTA logit-based methods, our method can achieve better performance in most settings. Furthermore, when combined with one of the SOTA feature-based methods ReviewKD, our method can achieve the best performance, showing that our method is compatible with feature-based methods.
Table 3. Top-1 and top-5 accuracy (%) on the ImageNet validation. We set ResNet-34 as the teacher and ResNet-18 as the student.
[…]
. distillation manner features logits Metric Teacher Student AT OFD CRD ReviewKD SimKD CAT-KD KD DKD DIST DHKD top-1 73.31 69.75 70.69 70.81 71.17 71.61 71.59 71.26 70.66 71.70 72.07 72.15 top-5 91.42 89.07 90.01 89.98 90.13 90.51 90.48 90.45 89.88 90.41 90.42 90.89
Table 4. Top-1 and top-5 accuracy (%) on the ImageNet validation. We set ResNet-50 as the teacher and MobileNet as the student.
[…]
. distillation manner features logits Metric Teacher Student AT OFD CRD ReviewKD SimKD CAT-KD KD DKD DIST DHKD top-1 76.16 68.87 69.56 71.25 71.37 72.56 72.25 72.24 68.58 72.05 73.24 72.99 top-5 92.86 88.76 89.33 90.34 90.41 91.00 90.86 91.13 88.98 91.05 91.12 91.45
ImageNet image classification.
The top-1 and top-5 validation accuracy on ImageNet is reported in Table 3 for cases where teachers and students share the same architecture and in Table 4 for cases where they have different architectures. We can find that our DHKD achieves comparable or even better performance than the existing methods. The success on the large-scale dataset further proves the effectiveness of our method.
4.2. Ablation Studies
@@
Table 5. Ablation studies on CIFAR-100. We choose one pair of models with the same architecture and another pair with different architectures. For the former pair, we set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student; for the latter pair, we set ResNet-50 as the teacher and MobileNet-V2 as the student.
[…]
. CE BinaryKL -Norm Dual Head Gradient Alignment Nonlinear Auxiliary Head resnet32$\times$4 resnet8$\times$4 ResNet-50 MBN-V2 72.50 64.60 75.15 68.52 N.A. N.A. 76.16 70.16 76.54 69.97 74.38 71.08
(a) Vanilla
(b) KD
(c) DHKD
(d) Teacher
Figure 5. t-SNE visualization of features learned by different methods. We do the visualization on the test set of CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
[…]
(a) Vanilla
(b) DKD
(c) DHKD Aux
(d) DHKD
Figure 6. The difference between the correlation matrices of the teacher’s and student’s logits. “DHKD Aux” is short for DHKD’s auxiliary head. The test set is CIFAR-100, with resnet32$\times$4 as the teacher and resnet8$\times$4 as the student. The performance sorting of four classifier heads is DHKD (76.54) $>$ DKD (76.37) $>$ DHKD’s Aux (76.20) $>$ Vanilla (72.17).
[…]
4.3. t-SNE Visualization of the Features
t-SNE (t-distributed Stochastic Neighbor Embedding) (Van der Maaten and Hinton, 2008) is a famous unsupervised non-linear dimensionality reduction technique for visualizing high-dimensional data. We use t-SNE to visualize the features (the outputs of the models’ backbones) learned by different models: vanilla student without distillation, models trained by KD, DHKD, and the teacher model.
4.4. Difference of Correlation Matrices
5. Conclusion
@@
6. Acknowledgement
The authors thank the anonymous reviewers for their careful reading of our manuscript and their many insightful comments and suggestions. This research is supported by the National Research Foundation, Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. This research is also supported by the Joint NTU-WeBank Research Centre on Fintech, Nanyang Technological University, Singapore. Sheng-Jun Huang is supported by the NSFC(U2441285, 62222605).
@@ Appendix B Supplementary Material about Neural Collapse
B.1. Detailed Definition of ETF and Neural Collapse
Definition 0 (Simplex Equiangular Tight Frame (Papyan et al., 2020)).
@@ B.1. Detailed Definition of ETF and Neural Collapse
(9) $\mathbf{M}=\sqrt{\frac{K}{K-1}}\mathbf{U}\left(\mathbf{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right),$
@@
(10) $\mathbf{m}_{i}^{\top}\mathbf{m}_{j}=\frac{K}{K-1}\delta_{i,j}-\frac{1}{K-1},\forall i,j\in[1,K],$
@@
B.2. Proof of Propositions
B.2.1. Proof of Proposition 1
@@
(11) $-\frac{\partial\mathcal{L}_{CE}}{\partial\bm{w}_{k}}=\underbrace{\sum_{i=1}^{n_{k}}\left(1-p_{k}\left(\bm{h}^{\mathcal{S}}_{k,i}\right)\right)\bm{h}^{\mathcal{S}}_{k,i}}_{\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}}+\underbrace{\left(-\sum_{k^{\prime}\neq k}^{K}\sum_{j=1}^{n_{k^{\prime}}}p_{k}\left(\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)}_{\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}},$
@@
(13) $-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL}}}{\partial\bm{w}_{k}}=$ $\underbrace{\sum_{i=1}^{n_{k}}\tau(q_{k}(\bm{h}^{\mathcal{T}}_{k,i})-q_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i}}_{\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}}+$
(14) $\underbrace{\left(-\sum_{k^{\prime}\neq k}^{K}\sum_{j=1}^{n_{k^{\prime}}}\tau(q_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j})-q_{k}(\bm{h}^{\mathcal{T}}_{k^{\prime},j}))\bm{h}^{\mathcal{S}}_{k^{\prime},j}\right)}_{\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}},$
@@
(15) $q_{k}(\bm{h})=\frac{1}{1+e^{-\bm{h}^{\top}\bm{w}_{k}/\tau}}.$
Based upon Equation 11, 13, we can prove Proposition 1.
B.2.2. Proof of Proposition 2
The definitions of $p_{k}(\bm{h})$ and $q_{k}(\bm{h})$ are the same as Equation 12 and 15.
@@ B.2.2. Proof of Proposition 2
(16) $-\frac{\partial\mathcal{L}_{CE}}{\partial\bm{h}}=\underbrace{(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}}+\underbrace{(-\sum\limits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k})}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}}.$
@@
(17) $-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL}}}{\partial\bm{w}_{k}}=$ $\underbrace{\tau(q_{c}(\bm{h}^{\mathcal{T}})-q_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}}+$
(18) $\underbrace{\left(-\tau\sum\limits_{k\neq c}^{K}(p_{k}(\bm{h}^{\mathcal{S}})-p_{k}(\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{k}\right)}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}}.$
Based upon Equation 16, 17, we can prove Proposition 2.
B.3. Propositions for the Modified Loss